OpenStudy (anonymous):

how do you find angle x in a right triangle that has angles sides A,B,C and leg adjacent is 19 and leg opposite is 22?

OpenStudy (anonymous):

ok, simple, jus use the SohCahToa principle. You already know the adjacent (a)and opsite (o), hence, you will have to use tangent. in SohCahToa, you see that Toa has Tan, Oposite and Adjacent. the formula is tanA = opposite/adjacent so, in your case it will be: tan(x) = 22/19 x = (tan^-1)22/19 x = 49.184 x= 49.2

OpenStudy (anonymous):

excellent :)